# Text clustering with LLM embeddings

### v3

@@ 3.5 Evaluation metrics
To comprehensively evaluate the quality of different embeddings and algorithm combinations, we used a diverse set of metrics. For external validation, since the original labels were available, we used the weighted $F1$-score (F1S) [36], the Adjusted Rand Index (ARI) [37], and the Homogeneity score (HS) [38]. The $F1$-score was computed to balance precision and recall in the presence of class imbalance. ARI was used to assess clustering outcomes while correcting for chance grouping, and HS was used to evaluate the degree to which each cluster is composed of data points primarily from one class. For internal validation, we employed the Silhouette Score (SS) [39] and the Calinski-Harabasz Index (CHI) [40], evaluating cluster coherence and separation without requiring ground truth. This multifaceted approach ensures a robust assessment, capturing both the alignment with known labels and the intrinsic structure of the generated clusters. These metrics collectively provide a balanced view of the performance, accounting for datasets with varying characteristics and sizes. The metrics and their corresponding formulas are presented in Table 4.
Table 4: Metrics used to assess clustering results, their type (external or internal), and their respective formulas. For $F1$-score (F1S), $C$ represents the number of classes, $w_{i}$ is the weight assigned to the $i$-th class, which is typically the proportion of that class within the dataset, and $F_{1,i}$ represents the $F1$-score computed for the $i$-th class. For the Adjusted Rand Index (ARI), $RI$ stands for Rand Index, $Expected\_RI$ refers to the expected value of the Rand Index under random label assignment (calculated using the contingency table marginals), and $Max\_RI$ is the maximum possible value of the Rand Index. For the Homogeneity Score (HS), $H(C|K)$ is the conditional entropy of the class distribution given the predicted cluster assignments and $H(C)$ is the entropy of the class distribution. For the Silhouette Score (SS), $N$ represents the total number of data points in the dataset, and $s(i)$ is the silhouette score for a single data point $i$, defined as $s(i)=\frac{b(i)-a(i)}{\max\{a(i),b(i)\}}$, where $a(i)$ is the average distance from the $i$-th data point to others points in the same cluster, $b(i)$ is the minimum mean distance from the $i$-th data point to points in a different cluster, minimised over all clusters. For the Calinski-Harabasz Index (CHI), $Tr(B_{k})$ is the trace of the between-group dispersion matrix that measures the between-cluster dispersion, $Tr(W_{k})$ is the trace of the within-cluster dispersion matrix which quantifies the within-cluster dispersion, $N$ refers to the number of data points, and $k$ indicates the number of clusters. The optimal value for all these metrics is the maximum.
@@ 3.6 Additional experiments
This section describes additional experiments in which we performed text summarisation (3.6.1) and tested LLM embeddings obtained from larger models (i.e., models with more parameters) prior to clustering (3.6.2). The purpose is to investigate whether such representations can improve the discriminability of features within text clusters.
@@ 3.6.1 Summarisation
This experiment aims to evaluate summarization as a tool for dimensionality reduction in text clustering by creating compact representations of the texts that encapsulate their semantic core without losing context. These experiments are hypothesised to streamline the clustering process, potentially leading to more coherent and interpretable clusters, even in large and complex datasets. This involves adding a summarization step after preprocessing and before the retrieval of embeddings. The models used in summarisation are described in Table 5. As an alternative approach to LLM-based models, we used the BERT-large-uncased summarisation model [5] implemented by the BERT summariser [41] to assess potential improvements in clustering achieved by utilising a lower-dimensionality model.
Table 5: Overview of text summarization models used in this study, each employing a transformer-based architecture. ‘Max tokens’ is the maximum input sequence length; BERT is a bidirectional encoder-only model trained on BookCorpus, a dataset consisting of 11,038 unpublished books and English Wikipedia (excluding lists, tables and headers); OpenAI is a GPT3-based decoder-only model built with multi-head attention blocks trained on extended dataset with reinforcement learning from human feedback; Falcon is a parameters-causal decoder-only model with multi-query attention mechanism trained on 1,500B tokens of RefinedWeb enhanced with curated corpora; LLaMA-2-chat is a LLaMA-2-based auto-regressive decoder-only model trained on 1, [^0] B tokens dataset collected by Meta with 2 trillion tokens, enriched with supervised fine-tuning, reinforcement learning from human feedback, and Ghost Attention mechanism.
@@
OpenAI gpt-3-5-turbo [42] 4096
Falcon falcon-7b [43] 2048
LLaMA-2-chat Llama-2-7b-chat-hf [44] 4096
@@
Table 6: Parameters used for summarisation with LLaMA-2 and Falcon models: $\mathsf{temperature}$ represents the value to modulate probabilities of the next token, $\mathsf{max\_length}$ defines the minimum length of the sequence to be generated, $\mathsf{do\_sample}$ is a parameter that determines whether to use sampling rather than greedy decoding, $\mathsf{top\_k}$ restricts the selection to the top $k$ tokens with the highest probabilities during top-$k$-filtering, and $\mathsf{num\_return\_sequences}$ is the number of independently computed returned sequences for each element in the batch.
@@ 3.6.2 Increasing model dimension
The original publications on LLMs underline the performance increase with larger model sizes in common sense topics, question answering, and code tasks [30, 31]. This experiment evaluates embeddings from various LLM sizes to analyse the impact of higher dimensionality models on the performance of clustering algorithms, aiming to determine if they enhance cluster cohesion and separation. For this purpose, we used embeddings obtained for models presented in Table 7.
@@
Table 7: Embeddings and corresponding models used in the dimensionality experiments. Here, size represents a number of parameters denoted in billions (bp), tokens refers to the embedding token size in trillions (trl).
@@ 4 Results and Discussion
Results demonstrate that OpenAI embeddings generally yield superior clustering performance on structured, formal texts compared to other methods based on most metrics (column “Total” in Table 8). The combination of the $k$-means algorithm and OpenAI’s embeddings yielded the highest values of ARI, $F1$-score, and HS in most experiments. This may be attributed to OpenAI’s embeddings being trained on a diverse array of Internet text, rendering them highly effective at capturing the diversity of language structures. Low values of SS and CHI for the same algorithm could indicate that, while clusters are homogeneous and aligned closely with the ground truth labels (suggesting good class separation and cluster purity), they may not be well-separated or compact as evaluated in a geometric space. This discrepancy can arise in cases where data has a high-dimensional or complex structure that external measures such as ARI, F1S, and HS capture effectively, but when projecting into a lower-dimensional space for SS and CHI, the clusters appear to overlap or vary widely in size, leading to a lower score for these spatial coherence metrics. Interestingly, for the case of the Reuters datasets (DS6), we observed that FuzzyCM consistently outperformed other clustering algorithms when paired with different embeddings. The potential reason behind this result may be related to the flexibility of fuzzy clustering, which allows data points to belong to multiple clusters with varying degrees of membership, making it particularly suitable for datasets with a large number of classes such as the Reuters dataset.
In the domain of open-source models, namely Falcon, LLaMA-2, and BERT, the latter emerged as the frontrunner. Given that BERT is designed to understand context and potentially due to the model’s lower dimensionality, these embeddings demonstrate good effectiveness in text clustering. In the comparative analysis of open-source LLM embeddings, Falcon-7b outperformed the LLaMA-2-7b across most datasets, demonstrating improved cluster quality and distinctiveness. This superiority may be attributed to Falcon-7b embeddings ability to better capture the salient linguistic features and semantic relationships within the texts since these embeddings were trained on a mixed corpus of text and code, as opposed to the LLaMA-2 embeddings, which are specialised for dialogues and Q&A contexts. Additionally, the CHI metric—measuring the dispersion ratio between and within clusters—is higher for Falcon-7b embeddings, suggesting that clusters are well separated and dense.
@@
In turn, results for the model size experiment, presented in Table 10, highlight the negative influence of number of parameters on clustering outcomes for Falcon, given that the larger model, Falcon-40b, produced lower ARI, $F1$-score, and HS for the CSTR and SyskillWebert datasets. This may be due to the fact that embeddings for the Falcon-40b model were created over a subset of the data used for the Falcon-7b embeddings (1.5 trillion tokens for Falcon-7b and 1 trillion tokens for Falcon-40b). In the case of LLaMA-2, models with sizes 7b and 13b were reportedly trained on identical datasets [44]. Results indicate that embeddings from the larger model, LLaMA-2-13b, outperform those from LLaMA-2-7b. This outcome is anticipated, as larger models with more parameters generally have a greater capacity to capture complex patterns and relationships in the data, leading to richer and more expressive embeddings.
Another perspective is given by Figure 1, which shows a visualisation of different LLM parameter sizes for the CSTR dataset using PCA and $t$-SNE. A noticeable artefact for LLaMA-2-7b (Figure 1a) and Falcon-40b (Figure 1d) is the lack of coherence in the four document classes, pointing to an unclear delineation in the feature space. Conversely, when employing LLaMA-2-13b (Figure 1b) and Falcon-7b (Figure 1c), classes have better separation, providing a possible explanation of the clustering results observed in Table 10. Increasing the number of tokens in Falcon yielded better results, aligning with existing literature that suggests models with larger vocabulary are more capable of capturing complex patterns in the data. Balancing computational requirements with clustering quality involves weighing the benefits of larger embeddings against their resource demands. As results show, larger number of parameters do not necessarily lead to improved clustering and require significantly more computational power and memory. Empirical evaluation is essential to determine if the improvements justify the additional costs, considering the specific task and resource constraints.
@@
Figure 1: Representation of different embeddings for the CSTR dataset, where PCA was used as preliminary dimensionality reduction algorithm and t-SNE for data projection into a lower-dimensional space.
In summary, these results provide valuable insights into the relationship between text embeddings, clustering algorithms, and dimensionality reduction techniques. We evaluated five embeddings and four clustering algorithms across five datasets, including the MN-DS dataset with hierarchical labels. Embeddings from the OpenAI model generally outperformed traditional techniques such as BERT and TF-IDF, aligning with previous studies. Conversely, LLaMA-2 and Falcon models produced inferior results compared to BERT across most metrics. The $k$-means algorithm was the most effective on most datasets, while FuzzyCM performed better on the Reuters dataset. Additionally, summarization models used for dimensionality reduction did not enhance clustering performance, and the use of higher-dimensional models displayed mixed results.
@@ 5 Limitations
Despite the insights revealed in this study, the available computational resources constrained the experiments, prohibiting us from conducting additional trials with summary generation and higher-dimensional models on more voluminous datasets. Consequently, the potential benefits of summaries in large-scale text clustering have yet to be thoroughly evaluated. Summarisation might behave differently when applied at scale or with distinct prompts, potentially providing more pronounced benefits or drawbacks depending on the complexity and diversity of the text data.
Furthermore, there is a possibility that embeddings computed for extra-large models like Falcon-180b or LLaMA-2-70b could deliver even substantial gains, but practical application is restrained by the significant computational demand that accompanies increased model size. This limitation potentially skews our understanding of the absolute efficacy of the embeddings computed for these models, as we could only infer performance improvements up to a specific size. These constraints prompt essential considerations for future research. Specifically, experiments with larger datasets and higher-dimension models would enable a more comprehensive and accurate understanding of the potentials and limitations of text clustering algorithms and their scalability in real-world applications.
@@ 6 Conclusions
In this study, we examined the impact that various embeddings—namely TF-IDF, BERT, OpenAI, LLaMA-2, and Falcon—and clustering algorithms have on grouping textual data. Through detailed exploration, we evaluated the efficacy of dimensionality reduction via summarisation and the role of model size on the clustering efficiency of various datasets. We found that OpenAI’s embeddings generally outperform other embeddings. BERT’s performance excelled amongst open-source alternatives, underscoring the potential of advanced models to positively influence text clustering results.
@@
These findings point towards continued research focused on developing strategies that leverage the strengths of advanced models while mitigating their computational demands. It is also critical to expand the scope of research to include more diverse text types, which will provide a more comprehensive understanding of clustering dynamics across different contexts and analyse embeddings of very recent or yet to be released models.

### v4

@@ 3.5 Evaluation Metrics
To comprehensively evaluate the quality of different embeddings and algorithm combinations, we used a diverse set of metrics. For external validation, since the original labels were available, we used the weighted $F1$-score (F1S) [33], the Adjusted Rand Index (ARI) [34], and the Homogeneity score (HS) [35]. The $F1$-score was computed to balance precision and recall in the presence of class imbalance. ARI was used to assess clustering outcomes while correcting for chance grouping, and HS was used to evaluate the degree to which each cluster is composed of data points primarily from one class. For internal validation, we employed the Silhouette Score (SS) [36] and the Calinski-Harabasz Index (CHI) [37], evaluating cluster coherence and separation without requiring ground truth. This multifaceted approach ensures a robust assessment, capturing both the alignment with known labels and the intrinsic structure of the generated clusters. These metrics collectively provide a balanced view of performance, accounting for datasets with varying characteristics and sizes. The metrics and their corresponding formulas are presented in Table 4.
Table 4: Metrics used to assess clustering results, their type (external or internal), and their respective formulas. For the $F1$-score (F1S), $C$ represents the number of classes, $w_{i}$ is the weight assigned to the $i$-th class, which is typically the proportion of that class within the dataset, and $F_{1,i}$ represents the $F1$-score computed for the $i$-th class. For the Adjusted Rand Index (ARI), $RI$ stands for Rand Index, $Expected\_RI$ refers to the expected value of the Rand Index under random label assignment (calculated using the contingency table marginals), and $Max\_RI$ is the maximum possible value of the Rand Index. For the Homogeneity Score (HS), $H(C|K)$ is the conditional entropy of the class distribution given the predicted cluster assignments, and $H(C)$ is the entropy of the class distribution. For the Silhouette Score (SS), $N$ represents the total number of data points in the dataset, and $s(i)$ is the silhouette score for a single data point $i$, defined as $s(i)=\frac{b(i)-a(i)}{\max\{a(i),b(i)\}}$, where $a(i)$ is the average distance from the $i$-th data point to other points in the same cluster, and $b(i)$ is the minimum mean distance from the $i$-th data point to points in a different cluster, minimised over all clusters. For the Calinski-Harabasz Index (CHI), $Tr(B_{k})$ is the trace of the between-group dispersion matrix that measures the between-cluster dispersion, $Tr(W_{k})$ is the trace of the within-cluster dispersion matrix, which quantifies the within-cluster dispersion, $N$ refers to the number of data points, and $k$ indicates the number of clusters. The optimal value for all these metrics is the maximum.
@@ 3.6 Additional Experiments
This section describes additional experiments in which we performed text summarisation (3.6.1) and tested LLM embeddings obtained from larger models (i.e., models with more parameters) prior to clustering (3.6.2). The purpose of these experiments is to investigate whether such representations can improve the discriminability of features within text clusters.
@@ 3.6.1 Summarisation
This experiment aims to evaluate summarisation as a tool for dimensionality reduction in text clustering by creating compact representations of the texts that encapsulate their semantic core without losing context. These experiments are hypothesised to streamline the clustering process, potentially leading to more coherent and interpretable clusters, even in large and complex datasets. This involves adding a summarisation step after preprocessing and before the retrieval of embeddings. The models used in summarisation are described in Table 5. As an alternative approach to LLM-based models, we used the BERT-large-uncased summarisation model [5] implemented by the BERT summariser [38] to assess potential improvements in clustering achieved by utilising a lower-dimensionality model.
Table 5: Overview of text summarisation models used in this study, each employing a transformer-based architecture. ‘Max Tokens’ is the maximum input sequence length. BERT is a bidirectional encoder-only model trained on BookCorpus, a dataset consisting of 11,038 unpublished books and English Wikipedia (excluding lists, tables, and headers); OpenAI is a GPT-3-based decoder-only model built with multi-head attention blocks, trained on an extended dataset with reinforcement learning from human feedback; Falcon is a causal decoder-only model with a multi-query attention mechanism trained on 1,500B tokens of RefinedWeb enhanced with curated corpora; LLaMA-2-chat is a LLaMA-2-based auto-regressive decoder-only model trained on a 1, [^0] B token dataset collected by Meta, enriched with supervised fine-tuning, reinforcement learning from human feedback, and the Ghost Attention mechanism.
@@
OpenAI gpt-3.5-turbo [39] 4096
Falcon falcon-7b [40, 27] 2048
LLaMA-2-chat Llama-2-7b-chat-hf [41, 28] 4096
@@
Table 6: Parameters used for summarisation with LLaMA-2 and Falcon models: $\mathsf{temperature}$ represents the value to modulate probabilities of the next token, $\mathsf{max\_length}$ defines the maximum length of the sequence to be generated, $\mathsf{do\_sample}$ is a parameter that determines whether to use sampling rather than greedy decoding, $\mathsf{top\_k}$ restricts the selection to the top $k$ tokens with the highest probabilities during top-$k$ filtering, and $\mathsf{num\_return\_sequences}$ is the number of independently computed returned sequences for each element in the batch.
@@ 3.6.2 Increasing Model Dimension
The original publications on LLMs underline the performance increase with larger model sizes in tasks such as common sense reasoning, question answering, and code tasks [27, 28]. This experiment evaluates embeddings from various LLM sizes to analyse the impact of higher-dimensional models on the performance of clustering algorithms, aiming to determine if they enhance cluster cohesion and separation. For this purpose, we used embeddings obtained from models presented in Table 7.
@@
Table 7: Embeddings and corresponding models used in the dimensionality experiments. Here, ‘Size’ represents the number of parameters denoted in billions (bp), and ‘Tokens’ refer to the embedding token size in trillions (trl).
@@ 4 Results and Discussion
Results demonstrate that OpenAI embeddings generally yield superior clustering performance on structured, formal texts compared to other methods based on most metrics (column ‘Total’ in Table 8). The combination of the $k$-means algorithm and OpenAI’s embeddings yielded the highest values of ARI, $F1$-score, and HS in most experiments. This may be attributed to OpenAI’s embeddings being trained on a diverse array of Internet text, rendering them highly effective at capturing the diversity of language structures.
Low values of SS and CHI for the same algorithm could indicate that, while clusters are homogeneous and aligned closely with the ground truth labels (suggesting good class separation and cluster purity), they may not be well-separated or compact as evaluated in a geometric space. This discrepancy can arise in cases where data has a high-dimensional or complex structure that external measures such as ARI, F1S, and HS capture effectively. However, when projecting into a lower-dimensional space for SS and CHI, the clusters appear to overlap or vary widely in size, leading to a lower score for these spatial coherence metrics. This highlights a challenge in text clustering: achieving both high cluster purity alongside tight cluster geometry.
Interestingly, for the case of the Reuters dataset (DS6), we observed that FuzzyCM consistently outperformed other clustering algorithms when paired with different embeddings. The potential reason behind this result may be related to the flexibility of fuzzy clustering, which allows data points to belong to multiple clusters with varying degrees of membership, making it particularly suitable for datasets with a large number of classes such as the Reuters dataset.
In the domain of open-source models, namely Falcon, LLaMA-2, and BERT, the latter emerged as the frontrunner. Given that BERT is designed to understand context and potentially due to the model’s lower dimensionality, these embeddings demonstrate good effectiveness in text clustering. In the comparative analysis of open-source LLM embeddings, Falcon-7b outperformed LLaMA-2-7b across most datasets, demonstrating improved cluster quality and distinctiveness. This superiority may be attributed to Falcon-7b embeddings’ ability to better capture salient linguistic features and semantic relationships within the texts since these embeddings were trained on a mixed corpus of text and code, as opposed to the LLaMA-2 embeddings, which are specialised for dialogues and Q&A contexts. Additionally, the CHI metric—measuring the dispersion ratio between and within clusters—is higher for Falcon-7b embeddings, suggesting that clusters are well-separated and dense.
@@
Results for the model size experiment, presented in Table 10, highlight the negative influence of the number of parameters on clustering outcomes for Falcon, given that the larger model, Falcon-40b, produced lower ARI, $F1$-score, and HS for the CSTR and SyskillWebert datasets. This may be due to the fact that embeddings for the Falcon-40b model were created over a subset of the data used for the Falcon-7b embeddings (1.5 trillion tokens for Falcon-7b and 1 trillion tokens for Falcon-40b). In the case of LLaMA-2, models with sizes 7b and 13b were reportedly trained on identical datasets [41]. Results indicate that embeddings from the larger model, LLaMA-2-13b, outperform those from LLaMA-2-7b. This outcome is anticipated, as larger models with more parameters generally have a greater capacity to capture complex patterns and relationships in the data, leading to richer and more expressive embeddings.
Another perspective is given by Figure 1, which shows a visualisation of different LLM parameter sizes for the CSTR dataset using PCA and $t$-SNE. A noticeable artefact for LLaMA-2-7b (Figure 1a) and Falcon-40b (Figure 1d) is the lack of coherence in the four document classes, indicating an unclear delimitation in the feature space. This observation is consistent with the results in Table 10, where these two models have the lowest F1S, ARI, and HS values. Conversely, when employing LLaMA-2-13b (Figure 1b) and Falcon-7b (Figure 1c), the classes exhibit better separation, which aligns with the clustering results observed in Table 10, where these two models achieved the highest values in 3 out of the 5 metrics calculated. Increasing the number of tokens in Falcon yielded better results, aligning with existing literature that suggests models trained on larger and more diverse corpora are more capable of capturing complex patterns in the data [43].
Balancing computational requirements with clustering quality involves weighing the benefits of larger embeddings against their resource demands. As results show, a larger number of parameters do not necessarily lead to improved clustering and require significantly more computational power and memory. Empirical evaluation is essential to determine if the improvements justify the additional costs, considering the specific task and resource constraints.
@@
Figure 1: Representation of different embeddings for the CSTR dataset, where PCA was used as a preliminary dimensionality reduction algorithm and t-SNE for data projection into a lower-dimensional space.
In summary, these results provide valuable insights into the relationship between text embeddings, clustering algorithms, and dimensionality reduction techniques. We evaluated five embeddings and four clustering algorithms across five datasets, including the MN-DS dataset with hierarchical labels. Embeddings from the OpenAI model generally outperformed traditional techniques such as BERT and TF-IDF, aligning with previous studies. Conversely, LLaMA-2 and Falcon models produced inferior results compared to BERT across most metrics. The $k$-means algorithm was the most effective on most datasets, while FuzzyCM performed better on the Reuters dataset. Additionally, summarisation models used for dimensionality reduction did not enhance clustering performance, and the use of higher-dimensional models displayed mixed results.
@@ 5 Limitations
Despite the insights revealed in this study, the available computational resources constrained the experiments. This limitation prohibited additional trials with summary generation, higher-dimensional models on more voluminous datasets, and fine-tuning of model parameters. Consequently, the potential benefits of summaries in large-scale text clustering have yet to be thoroughly evaluated. Summarisation might behave differently when applied at scale or with distinct prompts, potentially providing more pronounced benefits or drawbacks depending on the complexity and diversity of the text data.
Moreover, embeddings computed for extra-large models like Falcon-180b or LLaMA-2-70b could potentially deliver substantial gains. However, practical application is restrained by the significant computational demand that accompanies increased model size. This limitation potentially skews the understanding of the absolute efficacy of the embeddings computed for these models, as performance improvements could only be inferred up to a specific size.
Testing the selected clustering algorithms with a wider range of parameters could also provide additional insights into the results. However, due to the aforementioned computational limitations, a breadth-first approach was chosen—testing several different clustering algorithms—rather than a depth-first approach, which would involve experimenting with more parameters for each algorithm. For example, in AHC, the analysis was limited to using Euclidean distance with Ward linkage. Ward linkage was chosen because of its ability to minimize the variance within clusters, resulting in more compact and cohesive clusters, which is particularly beneficial for text data clustering. While this approach provides a comprehensive comparison across different methods, future research could explore parameter optimization in greater depth to further refine clustering outcomes.
These constraints prompt essential considerations for future research. Specifically, experiments with larger datasets and higher-dimension models would enable a more comprehensive and accurate understanding of the potentials and limitations of text clustering algorithms and their scalability in real-world applications.
@@ 6 Conclusions
In this study, we examined the impact that various embeddings—namely TF-IDF, BERT, OpenAI, LLaMA-2, and Falcon—and clustering algorithms have on grouping textual data. Through detailed exploration, we evaluated the efficacy of dimensionality reduction via summarisation and the role of model size on the clustering efficiency of various datasets. We found that OpenAI’s embeddings generally outperform other embeddings, with BERT’s performance excelling among open-source alternatives, underscoring the potential of advanced models to positively influence text clustering results.
@@
These findings point towards continued research focused on developing strategies that leverage the strengths of advanced models while mitigating their computational demands. It is also critical to expand the scope of research to include more diverse text types, which will provide a more comprehensive understanding of clustering dynamics across different contexts. Analyzing embeddings of very recent or yet-to-be-released models will also be important. Additionally, conducting parameter searches for clustering algorithms is essential to optimize their performance and ensure robust clustering outcomes across various datasets.